Multiply the following fractions:
(2/5) × 5 ¼ …........

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First convert the given mixed fraction into improper fraction.
= 5 whole 1⁄4
= 21/4
Now,
= (2/5) × (21/4)
By the rule Multiplication of fraction,
Product of fraction = (product of numerator)/ (product of denominator)
Then,
= (2 × 21)/ (5 × 4)
= (1 × 21)/ (5 × 2)
= (21/10)
= 2 whole 1/10

Thanks